OpenStudy (scott208):

A pizza is cut into six unequal slices (each cut starts at the center). The largest slice measures 90 degrees: If Larry eats the slices in order from the largest to the smallest, then the number of degrees spanned by a slice decreases at a constant rate. (So the second slice is smaller than the first by a certain number of degrees, then the third slice is smaller than the second slice by that same number of degrees, and so on.) What is the degree measure of the fifth slice Larry eats?

OpenStudy (avi13):

let decrease is x therefore (90)+(90-x)+(90-2x)........(90-5x)=360(sum of angle of centre of circle) solve for x and fitfh slice is 90-4x

OpenStudy (tackopp):

Hello!

OpenStudy (scott208):

so the second slice is 54 degrees

OpenStudy (scott208):

because 90-5x=360=54

OpenStudy (scott208):

so then for the third slice we do 90-4x=360?

OpenStudy (avi13):

no

OpenStudy (tackopp):

Okay. Since we have a pizza we know we are using a circle. The first slice is 90 degrees, which when subtracted form 360 leaves 270. Following the rule of needing a constant rate to be subtracted with each slice, we use 12 for our rate. Using 12, we find that the second slice measures 78, then 66, 54, 42, and finally 30 degrees of the last slice.

OpenStudy (scott208):

which one is wrong my answer for the second or third slice

OpenStudy (tackopp):

Second slice is ? and last slice is ?

OpenStudy (scott208):

thank you tackopp

OpenStudy (tackopp):

Your welcome!